A particle of mass 2 kg moves with an initial velocity of (4i+2j)m/s on the x-y plane . A force of F=2i-3j N acts on particle. The initial position of particle is (2m,3m). Then for y=3m
  • possible value of x is only x=2m
  • Possible value of x is not only x=2m but there exists some other value of x also
  • Time taken is 2s
  • All of these
provide me some explanation

Grade:11

1 Answers

Khimraj
6 years ago
Fx = 2i
Fy = -3j
ax = Fx/m = i
ay = Fy/m = -3j/2
ux = 4i
uy = 2j
x = 2 + (ux + axt)
y = 3 + ( uy + ayt)
for y=3
3 = 3 + (2 + -3t)
t = (2/3)s .
So at  t = 2/3 second particle wil again at y=3.
at this time position of x is
x = 2 +(4 + 2*(2/3)) = 22/3.
So correct option is Possible value of x is not only x=2m but there exists some other value of x also.
Hope it clears. If you like answer then please approve it.